3^(x+3)+3^x=84 (hint 3^x+3=3^x3^3) please help... Solve the equation
Found 2 solutions by Fombitz, jessica43:
Answer by Fombitz(13828)
Answer by jessica43(140)

The first step is to use the hint you were given (hint 3^(x+3)=(3^x)*(3^3)) to rewrite the problem:
(3^(x+3))+(3^x) = 84
((3^x)*(3^3))+(3^x) = 84
Now you can factor out the (3^x) and solve:
((3^x)*(3^3))+(3^x) = 84
(3^x)*((3^3)+1) = 84
(3^x)*(27+1) = 84
(3^x)*28 = 84
(3^x)= 3
x= 1
So your answer is x=1.